How long will it take to for your money to double

Given:

You have invested $1 today at 6% per year compounded yearly

Solve:

Approximately (to the nearest year) how many years will take for your money to double to $2 if invested at 6% per year compounded yearly?
Answer #1

Using the Rule of 72 the answer would be =72/6 =12 years,

The more accurate answer is calculated using the NPER function as follows:

=NPER(rate,pmt,pv,fv)

=NPER(6%,0,-1,2)

=11.90 or 12 years